Abigail gathered 11/40 as many apples as Ben. If Ben collected 116 more apples than Abigail, then how many apples does Abigail have?
Answer by stanbon(75887)

Abigail gathered 11/40 as many apples as Ben. If Ben collected 116 more apples than Abigail, then how many apples does Abigail have?
----------
Equations:
A = (11/40)B
B = A + 116
--------
Substitute for A and solve for "B"::
B = (11/40)B + 116
-----
(29/40)B = 116
B = (40/29)116
B = 160 (# of apples Ben has)
----
A = (11/40)B = (11/40)(160) = 11*4 = 44 (# of apples Abigail has.
------------
Cheers,
Stan H.